OpenStudy (anonymous):

Solve the following equation! 10e^x-45e^-x-15=0 Thanks in Advance!

OpenStudy (anonymous):

multiply by \(e^x\) and then solve a quadratic equation in \(e^x\)

OpenStudy (anonymous):

Could you please show how? :L

OpenStudy (anonymous):

\[10e^x-45e^{-x}-15=0\] multiply by \(e^x\) and get \[10e^{2x}-45-15e^x=0\]

OpenStudy (anonymous):

divide by 5 gives \[2e^{2x}-3e^x-9=0\] this is a quadratic equation in \(e^x\) as it is the same as \[2(e^x)^2-3e^x-9=0\]

OpenStudy (anonymous):

or you can think of it as letting \(z=e^x\) and you get \[2z^2-3z-9=0\] if you are lucky it factors

OpenStudy (anonymous):

Okay man thank you so much! I never thought of doing it that way!

OpenStudy (anonymous):

can you factor this one? don't forget, that when you solve for \(z\) you then have to write \(z=e^x\) and solve for \(x\), which only requires taking the log

OpenStudy (anonymous):

Yeah, or ln, I just needed to get it to where it's a quadratic equation.

OpenStudy (anonymous):

oh and throw out the negative solution one solution is \(z=-\frac{3}{2}\) but \(e^x=-\frac{3}{2}\) is not possible

OpenStudy (anonymous):

Yeah that's what I did. There's only 1 solution.

OpenStudy (anonymous):

k good!

OpenStudy (anonymous):

Thanks alot! You're awesome!